The town of sprinfield purchased a Chevrolet, Ford and Toyota for total of $ 66,000. The Ford was $ 2000 more than the Cheverlot and the Toyota was $ 2000 more than the Ford. What is the price of each car?
Kindly help me. Thank you.

Answer by ptaylor(2198)

Let x=cost of Chevrolet
Then (x+2000)=cost of Ford
And (x+2000)+2000=cost of Toyota
Now we are told that the above three prices add up to 66,000 dollars. So:
x+(x+2000)+((x+2000)+2000)=66,000 get rid of parens (actually we didn't need them in the first place)
x+x+2000+x+4000=66,000 start collecting like terms
3x+6000=66,000 subtract 6000 from both sides
3x+6000-6000=66,000-6000 finish collecting like terms
3x=60,000 divide both sides by 3
x=20,000--------------------------cost of Chevrolet
x+2000=20,000+2,000=22,000 ----------------cost of Ford
x+4000=20,000+4,000=24,000-----------------------Cost of Toyota
CK
20,000+22,000+24,000=66,000
66,000=66,000
Hope this helps---ptaylor